A process filling small bottles with baby formula has a target of 3.2 ouncesplus or minus0.270 ounce. Two hundred bottles from the process were sampled. The results showed the average amount of formula placed in the bottles to be 3.150 ounces. The standard deviation of the amounts was 0.072 ounce. Determine the value of Upper C Subscript pk. Roughly what proportion of bottles meet the​ specifications?
Answer #1

USL - upper specification limit = 3.2 +0.270 = 3.47 ISL z lower specification limit=3.2 -0. 270 = 2.93 mean = n=3.5 standard